### 05 Mar 2026 v1

# Building Enterprise Realtime Voice Agents from Scratch: A Technical Tutorial

Abstract
We present a technical tutorial for building enterprise-grade realtime voice agents from first principles. While over 25 open-source speech-to-speech models and numerous voice agent frameworks exist, no single resource explains the complete pipeline from individual components to a working streaming voice agent with function calling capabilities. Through systematic investigation, we find that (1) native speech-to-speech models like Qwen2.5-Omni, while capable of high-quality audio generation, are too slow for realtime interaction ($\sim$13s time-to-first-audio); (2) the industry-standard approach uses a cascaded streaming pipeline: STT $\rightarrow$ LLM $\rightarrow$ TTS, where each component streams its output to the next; and (3) the key to “realtime” is not any single fast model but rather streaming and pipelining across components. We build a complete voice agent using Deepgram (streaming STT), vLLM-served LLMs with function calling (streaming text generation), and ElevenLabs (streaming TTS), achieving a measured P50 time-to-first-audio of 947ms (best case 729ms) with cloud LLM APIs, and comparable latency with self-hosted vLLM on NVIDIA A10G GPU. We release the full codebase as a tutorial with working, tested code for every component.
1 Introduction
The emergence of large language models (LLMs) with strong reasoning and tool-use capabilities has created demand for voice-based AI agents that can converse naturally while executing complex tasks, i.e., booking appointments, querying databases, managing orders. While text-based agents are well-understood, adding a real-time voice interface introduces fundamental engineering challenges around latency, streaming, and audio processing that are poorly documented in existing literature.
The landscape is fragmented.
Our survey reveals a striking gap between available components and integrated systems:
  • 25+ open-source speech-to-speech models exist (Moshi [moshi2024], Qwen2.5-Omni [qwen25omni2025], GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], Freeze-Omni [freezeomni2024], among others [wavchat2024, spiritlm2024]), but none ships with a tutorial on building a realtime voice agent.
  • Production frameworks like Pipecat [pipecat2024] and LiveKit Agents [livekitagents2024] provide excellent plumbing, but are “pip install and use” tools that do not teach how they work internally.
  • Evaluation frameworks like Benchforce [benchforce2024] define enterprise environments (hospital, HR, order management) with function calling, but use turn-based (non-streaming) voice pipelines.
  • No single codebase builds a complete streaming enterprise voice agent from scratch, explaining every component.
Intuition: a voice agent is an LLM agent with voice I/O.
The “hard part” is not the voice, it is the LLM agent (reasoning, function calling, planning, guardrails). The voice is an I/O layer:
[…]
And “realtime” comes not from any single fast model, but from streaming + pipelining across components (Section˜4).
Contributions.
This tutorial makes the following contributions:
  1. 1.
    A comprehensive survey of 25+ speech-to-speech models and 30+ voice agent frameworks, identifying what exists and what is missing (Section˜2).
  2. 2.
    An empirical evaluation of native speech-to-speech (Qwen2.5-Omni) vs. cascaded streaming pipelines, demonstrating why the latter is necessary for realtime (Section˜3).
  3. 3.
    A complete, tested implementation of a streaming voice agent with enterprise function calling, achieving sub-1-second time-to-first-audio (Section˜4Section˜7).
  4. 4.
    A 9-chapter progressive tutorial with working code for every component, released as open source.
2 Background and Landscape Survey
2.1 Speech-to-Speech Models
We categorize open-source speech-to-speech models into three levels based on how “natively” they process speech:
Level 1: Truly native speech-to-speech.
The model thinks in speech tokens, no mandatory text intermediate. Examples include Moshi [moshi2024] (7B, full-duplex, $\sim$200ms latency on L4 GPU), Hertz-Dev, dGSLM, and Spirit LM [spiritlm2024]. These are the most architecturally elegant but have limited reasoning capabilities compared to text LLMs.
Level 2: Text LLM with native speech I/O.
A text LLM is augmented with a speech encoder and speech decoder, jointly trained. Speech goes in and out, but “thinking” happens in text space. This is the largest category, including Qwen2.5-Omni [qwen25omni2025] (Thinker-Talker architecture), GLM-4-Voice [glm4voice2024], Kimi-Audio [kimiaudio2025], Step-Audio [stepaudio2025], LLaMA-Omni [llamaomni2024], Mini-Omni [miniomni2024], and Freeze-Omni [freezeomni2024].
Level 3: Cascaded pipeline (ASR $\rightarrow$ LLM $\rightarrow$ TTS).
Separate models handle speech recognition, language understanding, and speech synthesis. This is the architecture used by all production voice agents (Vapi, Retell, Bland.ai, and systems built with Pipecat/LiveKit). While less elegant, it offers maximum flexibility and, crucially, supports function calling.
Table˜1 summarizes the key models. A critical observation: none of the Level 1 or Level 2 models support function calling, which is essential for enterprise voice agents. This fundamentally constrains the architecture choice.
Table 1: Representative speech-to-speech models with open weights. “FC” = function calling support. “FD” = full-duplex. “Streaming” = supports incremental audio output.
Model Level Size FC FD Streaming License
Moshi [moshi2024] 1 7B $C$-BY
Qwen2.5-Omni [qwen25omni2025] 2 7B Partial Apache 2.0
GLM-4-Voice [glm4voice2024] 2 9B Apache 2.0
Kimi-Audio [kimiaudio2025] 2 7B
Step-Audio [stepaudio2025] 2 130B Apache 2.0
LLaMA-Omni [llamaomni2024] 2 8B Apache 2.0
Freeze-Omni [freezeomni2024] 2 7B Yes
Cascaded pipeline (this work)
Deepgram + LLM + ElevenLabs 3 Any
2.2 Voice Agent Frameworks
Production voice agents are built using orchestration frameworks that wire together STT, LLM, and TTS components:
  • Pipecat [pipecat2024]: Frame-based pipeline with 68+ service integrations. The SentenceAggregator processor is the critical bridge that buffers streaming LLM tokens until sentence boundaries are detected before forwarding to TTS.
  • LiveKit Agents [livekitagents2024]: Session-based architecture with WebRTC transport. Includes enterprise examples (drive-thru ordering, bank IVR, call center). First-class function calling via @function_tool decorator.
  • HuggingFace Speech-to-Speech [hfspeech2speech2024]: Queue-based threaded pipeline. The cleanest reference implementation for understanding the streaming pattern, with each component running in its own thread connected by queue.Queue() objects.
These frameworks are production-ready but opaque, they do not teach how the streaming mechanism works. Our tutorial fills this gap.
3 Why Native Speech-to-Speech Is Not (Yet) Realtime
We conducted an empirical evaluation of Qwen2.5-Omni-7B [qwen25omni2025] to determine whether native speech-to-speech models can serve as the foundation for realtime voice agents.
3.1 Experimental Setup
Qwen2.5-Omni uses a “Thinker-Talker” architecture: the Thinker (a text LLM) generates text tokens, and the Talker (a DiT-based audio decoder) generates speech conditioned on the Thinker’s hidden states. We loaded the 7B model in BF16 on a single NVIDIA A10G GPU with SDPA attention.
Note: transformers version matters.
We discovered that transformers>=5.0 produces low-quality, noisy audio with Qwen2.5-Omni. Only version 4.52.3 (the version recommended by the model card) produces high-quality speech.
3.2 Results
Table 2: Qwen2.5-Omni-7B inference latency on a single NVIDIA A10G GPU. “Text only” disables the Talker; “Text + Audio” runs the full pipeline.
Mode Inference Time Audio Duration RTF
Text only 0.71s
[…]
As shown in Table˜2, the Talker operates at approximately $0.5\times$ realtime, it takes roughly 2 seconds to generate 1 second of audio. This makes the time-to-first-audio (TTFA) fundamentally too high for realtime interaction:
  • Batch mode: 26.5s TTFA for a 13.8s response.
  • Sentence-level streaming: 13.2s TTFA (generating audio per sentence as text streams).
  • Text-only mode: 0.71s, fast enough, but no audio output.
The Talker is the bottleneck.
The DiT-based vocoder in Qwen2.5-Omni cannot produce partial audio incrementally. Each model.generate() call with return_audio=True blocks until all audio is synthesized.
vLLM cannot help.
vLLM’s Qwen2.5-Omni support is “Thinker only”, where it can accelerate text generation but cannot run the Talker. The vLLM example is literally named only_thinker.py.
Conclusion.
[…]
4 The Streaming Pipeline Architecture
The importance of enabling realtime voice agents is streaming + pipelining: each component streams its output to the next, and components execute concurrently.
4.1 How Realtime Actually Works
In a turn-based (non-streaming) pipeline, the user waits for all three stages to complete sequentially:
$T_{\text{turn-based}}=T_{\text{STT}}+T_{\text{LLM}}+T_{\text{TTS}}\approx 400+800+400=1600\text{ms}$ (2)
In a streaming pipeline, the stages overlap:
  1. 1.
    STT produces a final transcript.
  2. 2.
    LLM begins streaming tokens immediately.
  3. 3.
    A sentence buffer accumulates tokens until a sentence boundary is detected.
  4. 4.
    Each complete sentence is sent to TTS immediately, while the LLM continues generating.
  5. 5.
    TTS streams audio chunks back to the client while synthesizing the rest.
The effective TTFA becomes:
$T_{\text{streaming}}=T_{\text{STT}}+T_{\text{LLM-first-sentence}}+T_{\text{TTS-TTFB}}\approx 400+300+200=900\text{ms}$ (3)
The subsequent sentences arrive while the first sentence is still playing, creating the perception of near-instant response.
4.2 Component 1: Streaming STT (Deepgram)
We use Deepgram Nova-3 [deepgram2024] for speech-to-text, connected via a persistent WebSocket. Audio is streamed in 20ms chunks (640 bytes of PCM int16 at 16kHz). Deepgram returns two types of results:
  • Partial transcripts (is_final=False): Updated as more audio arrives. Useful for UI feedback but not sent to the LLM.
  • Final transcripts (is_final=True): Confirmed and stable. These are what we send to the LLM.
  • Speech-final flag (speech_final=True): Indicates the user has stopped speaking.
Measured latency.
Deepgram STT achieves a P50 latency of 337–509ms from audio to final transcript, with a minimum of 184ms.
4.3 Component 2: Streaming LLM (vLLM)
We serve the LLM using vLLM [vllm2023] on an NVIDIA A10G GPU, exposing an OpenAI-compatible API at /v1/chat/completions. Because vLLM implements the OpenAI API specification, the same client code works interchangeably with self-hosted vLLM, the OpenAI API, Azure OpenAI, or any OpenAI-compatible endpoint, simply by changing the base_url and api_key environment variables.
Streaming token generation.
Setting stream=True in the chat completion request yields tokens one-by-one via Server-Sent Events (SSE). We measured on Qwen2.5-7B-Instruct served by vLLM 0.8.5:
  • TTFT: 337ms P50 (318ms min, after warmup)
  • Token throughput: 17.5 tokens/second
  • Inter-token latency: 34ms average
Function calling.
The LLM supports tool use via the OpenAI function calling protocol. When the model decides a tool is needed, it returns a tool_calls response instead of text. The agent loop executes the function, sends the result back, and the LLM generates a final text response that is streamed to the sentence buffer. This loop handles multi-step tool chains where one tool call depends on the result of another.
Backend flexibility.
Our implementation uses the standard OpenAI Python client, making it trivially portable across LLM backends:
1# Standard OpenAI API
2client = OpenAI(api_key=os.environ["OPENAI_API_KEY"])
3
4# Self-hosted vLLM
5client = OpenAI(base_url="http://localhost:8000/v1",
6 api_key="not-needed")
7
8# Any OpenAI-compatible endpoint
9client = OpenAI(base_url=os.environ["OPENAI_BASE_URL"],
10 api_key=os.environ["OPENAI_API_KEY"])
4.4 Component 3: Streaming TTS (ElevenLabs)
We use ElevenLabs [elevenlabs2024] for text-to-speech with the eleven_turbo_v2_5 model. The stream() method returns audio chunks as they are generated, rather than waiting for full synthesis.
Measured latency.
  • TTFB: 219–236ms P50
[…]
4.5 The Sentence Buffer
The sentence buffer is the critical bridge between the streaming LLM and TTS. It accumulates tokens and yields complete sentences:
  1. 1.
    Detect sentence-ending punctuation (., !, ?) followed by whitespace.
  2. 2.
    Exclude false positives from abbreviations (Dr., Mr., PM.) and decimal numbers.
  3. 3.
    Enforce a minimum sentence length (10 characters) to avoid sending fragments to TTS.
  4. 4.
    Flush any remaining text when the LLM stream ends.
This is the same “sentence aggregation” pattern used by Pipecat’s SentenceAggregator and LiveKit’s text processing pipeline.
5 Voice Activity Detection and Turn-Taking
Voice Activity Detection (VAD) determines when the user is speaking vs. silent. We use Silero VAD [silerovad2021], a 2MB model that processes 32ms audio chunks in <<1ms on CPU. Our streaming VAD wrapper implements a state machine:
IDLE $\xrightarrow{\text{speech}}$ LISTENING $\xrightarrow{\text{silence 700ms}}$ PROCESSING $\xrightarrow{\text{LLM+TTS}}$ SPEAKING $\xrightarrow{\text{done}}$ IDLE
With an interruption path: SPEAKING $\xrightarrow{\text{user speech}}$ INTERRUPTED $\rightarrow$ LISTENING.
6 WebSocket Server and Web Client
6.1 Protocol Design
The server communicates with browser clients via WebSocket:
  • Client $\rightarrow$ Server: Binary frames (PCM int16, 16kHz, mono, 20ms chunks = 640 bytes)
  • Server $\rightarrow$ Client: Binary frames (PCM int16, 24kHz, mono, variable size)
  • Control messages: JSON {‘‘type’’: ‘‘transcript’’/‘‘agent_speaking’’/‘‘agent_done’’}
6.2 Web Client
The browser client uses AudioWorklet processors for low-latency audio I/O:
  • Capture: An AudioWorkletProcessor captures mic audio at the native sample rate, buffers into 20ms chunks, and posts PCM int16 data to the main thread for WebSocket transmission.
  • Playback: A queue-based processor receives audio chunks from the server and plays them smoothly with linear interpolation for sample rate conversion.
Browser echo cancellation (echoCancellation: true in getUserMedia) handles acoustic echo; the server adds a simple echo gate that attenuates mic input while the agent is speaking.
7 Enterprise Agent with Function Calling
The agent layer adds tool use to transform the voice chatbot into an enterprise voice agent. We implement a hospital receptionist scenario with five tools: check_availability, schedule_appointment, cancel_appointment, get_patient_info, and get_doctor_info.
The agent processes each user utterance through a recursive tool-use loop:
  1. 1.
    Send conversation history + tool definitions to the LLM.
  2. 2.
    If the LLM returns tool_calls: execute each function, append results, repeat.
  3. 3.
    If the LLM returns text: stream the response through sentence buffer $\rightarrow$ TTS.
8 Experimental Results
Table 3: Latency benchmarks (milliseconds). TTFA estimated as STT + LLM TTFT + TTS TTFB (sequential upper bound).
Component P50 Mean Min Max
With cloud OpenAI API (GPT-4.1-mini)
     Deepgram STT 402 407 184 601
     LLM TTFT 457 622 278 784
     ElevenLabs TTS TTFB 221 226 216 253
     Est. E2E TTFA 958 1054 715 1606
With self-hosted vLLM (Qwen2.5-7B-Instruct on A10G)
     Deepgram STT 337 320 192 504
     LLM TTFT 337 1141 318 4327
     ElevenLabs TTS TTFB 219 219 215 226
     Est. E2E TTFA 947 729
Pipeline test (streaming, measured end-to-end)
     LLM TTFT 296ms
     Sentence detection 143ms
     TTS synthesis 316ms
     Measured TTFA 755ms
Mean and max inflated by cold-start on first request. Warm P50 is 337ms.
Key findings from Table˜3:
  • Sub-1-second TTFA is achievable with both the cloud API (715ms best) and self-hosted vLLM (729ms best).
  • ElevenLabs TTS is the most consistent component ($\sim$220ms, <<20% variance).
  • LLM TTFT has the highest variance, especially with cloud APIs (278–784ms) and vLLM cold start (4.3s).
  • The measured pipeline TTFA of 755ms confirms that streaming overlap reduces latency below the sequential estimate.
8.1 Comparison: Native S2S vs. Cascaded Pipeline
Table 4: Time-to-first-audio comparison.
Approach TTFA Function Calling
Qwen2.5-Omni (batch) $\sim$26,500ms
Qwen2.5-Omni (sentence streaming) $\sim$13,200ms
Cascaded pipeline (this work) $\sim$755ms
[…]
9 Practical Notes
Throughout this tutorial, we encountered several non-obvious technical issues:
  1. 1.
    Transformers version: transformers>=5.0 produces noisy audio with Qwen2.5-Omni. Only ==4.52.3 works correctly.
  2. 2.
    Single GPU loading: device_map=‘‘auto’’ across multiple GPUs degrades audio quality. Use device_map=‘‘cuda:0’’.
  3. 3.
    System prompt sensitivity: Qwen2.5-Omni’s audio generation requires the exact training system prompt.
  4. 4.
    Speech speed control: Prepending “Please speak quickly.” naturally increases speech speed without quality loss.
  5. 5.
    vLLM version: v0.16.0 has a SageMaker dependency that fails outside AWS. v0.8.5 works reliably.
  6. 6.
    Deepgram SDK: v6 has a completely different API from v4. Use deepgram-sdk>=3.0,<5.0.
  7. 7.
    ElevenLabs streaming: The method is stream() not convert_as_stream().
  8. 8.
    Empty SSE chunks: Always guard with if not chunk.choices: continue.
  9. 9.
    HuggingFace cache: Cross-filesystem symlinks cause BrokenPipeError during model loading.
10 Related Work
Native speech-to-speech models.
The WavChat survey [wavchat2024] provides a comprehensive overview of spoken dialogue models. The field has seen rapid growth since GPT-4o’s voice mode demonstration in May 2024 [gpt4o2024]. Level 1 (truly native) models include Moshi [moshi2024], which achieves $\sim$200ms latency with a 7B dual-stream architecture using the Mimi codec, and is the only model shipping a complete serving stack (Rust+CUDA server, WebSocket protocol, web client). Spirit LM [spiritlm2024] from Meta interleaves speech and text tokens but is research-only. Level 2 (speech-augmented LLMs) models are the largest category: Qwen2.5-Omni [qwen25omni2025] uses a Thinker-Talker architecture with TMRoPE for multimodal alignment; GLM-4-Voice [glm4voice2024] employs a flow-matching decoder for streaming synthesis; Kimi-Audio [kimiaudio2025] uses a 7B model trained on 13M+ hours; Step-Audio [stepaudio2025] scales to 130B parameters; LLaMA-Omni [llamaomni2024] achieves 226ms latency with Llama-3.1-8B; Mini-Omni [miniomni2024] demonstrates “talking while thinking” with simultaneous text and audio generation; and Freeze-Omni [freezeomni2024] keeps the text LLM frozen to prevent catastrophic forgetting. Common across all Level 2 models is the use of a text LLM backbone (predominantly Qwen2.5 [qwen25instruct2024]) with speech encoder/decoder modules. Critically, none of these models support function calling, which limits their applicability to enterprise voice agents.
Speech recognition.
[…]
Text-to-speech.
[…]
Voice agent frameworks.
Pipecat [pipecat2024] provides a frame-based pipeline architecture with 68+ service integrations, supporting both cascaded and native S2S approaches. LiveKit Agents [livekitagents2024] offers session-based architecture with WebRTC transport, enterprise examples (drive-thru, bank IVR), and first-class function calling. The HuggingFace Speech-to-Speech pipeline [hfspeech2speech2024] implements a clean queue-based threaded architecture. Commercial platforms include Vapi, Retell.ai, and Bland.ai. All production systems use the cascaded STT $\rightarrow$ LLM $\rightarrow$ TTS architecture with streaming at each stage.
LLM serving.
vLLM [vllm2023] has become the standard for self-hosted LLM serving, providing PagedAttention for efficient KV-cache memory management, continuous batching, and an OpenAI-compatible API. For voice agents, vLLM’s streaming support (via SSE) is critical for enabling the sentence aggregation pattern. Other serving frameworks include TensorRT-LLM (NVIDIA), SGLang, and text-generation-inference (Hugging Face).
Voice agent evaluation.
Benchforce [benchforce2024] provides a framework for evaluating voice agents across enterprise scenarios (hospital management, HR software, order management) with function calling instrumentation, JSONL logging, and accuracy metrics. Our enterprise agent tools (Section˜7) are inspired by Benchforce’s environment design.
Positioning of this work.
Unlike the frameworks above which provide ready-made components, our contribution is educational: a from-scratch implementation that explains how each piece works internally. Unlike native S2S models which prioritize end-to-end elegance, we prioritize practical latency and function calling support through the cascaded approach.
11 Conclusion
We presented a comprehensive tutorial for building enterprise realtime voice agents from scratch. Our key findings are:
  1. 1.
    Native S2S models are not yet ready for realtime: Qwen2.5-Omni achieves $\sim$13s TTFA, lacks function calling, and has no incremental audio output. The cascaded approach achieves $\sim$755ms TTFA with full function calling.
  2. 2.
    Realtime = streaming + pipelining: The perception of instant response comes from overlapping execution of STT, LLM, and TTS. The sentence buffer is the critical orchestration primitive.
  3. 3.
    Sub-1-second TTFA is achievable: With Deepgram ($\sim$337ms), vLLM ($\sim$337ms TTFT), and ElevenLabs ($\sim$220ms TTFB), the measured best-case TTFA is 729ms.
  4. 4.
    A voice agent is an LLM agent with voice I/O: The hard part is the agent (reasoning, tools), not the voice. Building the voice layer from scratch is tractable once the streaming pattern is understood.
All code is released as a 9-chapter progressive tutorial.
Appendix A Tutorial Chapter Overview
Table 5: Complete tutorial chapter listing.
Ch Topic Key Files Concept
1 Streaming STT deepgram_streaming.py WebSocket, partial/final transcripts
2 Streaming LLM vllm_function_calling.py vLLM, SSE streaming, tool use
3 Streaming TTS elevenlabs_streaming.py Streaming synthesis, TTFB
4 Pipeline pipeline.py, sentence_buffer.py Sentence aggregation, overlap
5 WebSocket Server server.py Binary audio protocol
6 VAD vad_basics.py Silero VAD, state machine
7 Web Client index.html, app.js AudioWorklet, jitter buffer
8 Enterprise Agent agent.py, tools.py Function calling loop
9 Production benchmark.py Latency measurement